### v3

# Harm Amplification in Text-to-Image Models

Abstract
Warning: The content of this paper as well as some blurred images shown include references to nudity, sexualization, violence, and gore.
Text-to-image (T2I) models have emerged as a significant advancement in generative AI; however, there exist safety concerns regarding their potential to produce harmful image outputs even when users input seemingly safe prompts. This phenomenon, where T2I models generate harmful representations that were not explicit in the input prompt, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms. Our paper addresses this issue by formalizing a definition for this phenomenon which we term harm amplification. We further contribute to the field by developing a framework of methodologies to quantify harm amplification in which we consider the harm of the model output in the context of user input. We then empirically examine how to apply these different methodologies to simulate real-world deployment scenarios including a quantification of disparate impacts across genders resulting from harm amplification. Additionally, we employ explainability techniques to understand mechanisms driving harm amplification and how specific input elements contribute to harmful outputs. Together, our work aims to offer researchers tools to comprehensively address safety challenges in T2I systems and contribute to the responsible deployment of generative AI models.
Introduction
Generative text-to-image (T2I) systems allow users to create new image content in response to a text input prompt. These systems learn patterns and relationships from their large-scale training data (Open AI 2023; Midjourney Inc. 2022; Stability AI 2022). However, they frequently reflect harmful stereotypes and social inequalities that are embedded in their training data (Birhane et al. 2023; Birhane, Prabhu, and Kahembwe 2021; Kirk et al. 2021; Wang et al. 2022), which are subsequently reproduced in generated imagery (Cho, Zala, and Bansal 2023). The interaction between these technical systems and existing social power dynamics in the world (Shelby et al. 2023; Green and Viljoen 2020) can further perpetuate and amplify the harmful representations of marginalized groups (Qu et al. 2023; Qadri et al. 2023; Rando et al. 2022b).
Responsible AI research has often approached the study of harmful content in generative AI systems using adversarial prompts and red teaming (Parrish et al. 2023; Rando et al. 2022b). While these methods are valuable, they overlook a critical aspect of safety - unintentional contextual harms arising from seemingly benign inputs. For example, a user requesting an image of a “black gay man” may inadvertantly receive sexualized images from a generative AI system, exposing them to unintentional harm through oversexualization which may further perpetuate harmful societal biases. Even with safety filters in place (Hao et al. 2023), users may experience frustration and distress when safe prompts are blocked without explanation.
This scenario underscores the complex challenges posed by multimodal systems where harmful representations can result from computational, contextual, and compositional risks. While responsible AI scholarship acknowledges the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities” (Shelby et al. 2023, p. 733), there is currently no precise definition of multimodal safety harm amplification.
To address these challenges, we provide the first formalization of harm amplification and make the following research contributions:
  • A formal definition of harm amplification for T2I models as occurring when the generated image from a T2I model reflects harmful or unsafe representations that were not explicit in the text input.
  • A framework that includes three methods for quantifying and measuring harm amplification instances in T2I models.
  • An empirical examination of patterns in harm amplification including quantifying its disparate impact across genders and applying explainability techniques to identify aspects of the user input that contribute to harmful output generation.
In what follows, we first orient our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification; followed by a description of our novel methodologies and their efficacy in measuring harm amplification. We then examine patterns of harm amplification within our data and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems, arguing for greater attention to harm amplification in responsible AI research and practice.
Related Work
This research builds on existing responsible AI literature addressing harm, safety, and multimodal generative AI, specifically T2I systems. Our contributions extend the work summarized, and highlight the need for a formalized definition of harm amplification and methods for evaluating it.
Sociotechnical Safety in T2I Systems and Harm Reduction
AI safety encompass a range of diverse sociotechnical issues that impact the well-being of people and the environment. A systems approach to AI safety identifies hazards (Dobbe 2022) and integrates safety considerations into the development process (Leveson 2016) to minimize potential harms (Dobbe 2022). Addressing social and ethical safety in T2I systems requires examining the interactions between AI components and social dynamics (Rismani et al. 2023; Weidinger et al. 2023). AI systems can replicate social dynamics (Benjamin 2019) or even amplify harmful representations that perpetuate societal biases (Qadri et al. 2023; Bianchi et al. 2023).
An effective governance approach to cultivating “safe” AI systems is to develop clear safety requirements (Leveson 2020), with attention to the sociotechnical nature of AI systems (Dobbe 2022) and potential harms (Shelby et al. 2023). There is increasing recognition of representational harms in T2I systems (Cho, Zala, and Bansal 2023; Díaz et al. 2023; Bird, Ungless, and Kasirzadeh 2023), which are the socially constructed beliefs about different social groups that reinforce unjust hierarchies (Wang et al. 2022; Barocas et al. 2017; Barocas, Hardt, and Narayanan 2019), and manifest in different ways (Katzman et al. 2021). Representational harms are sociotechnical in nature (Shelby et al. 2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations (Birhane, Prabhu, and Kahembwe 2021).
Gender inequalities, exacerbated by AI technologies, illustrate a significant global challenge (Lütz 2023) and exemplify a specific instance of representational harms. Gender, a social construct, shapes expectations around masculinity, and femininity, and beyond (Lorber 1996; West and Zimmerman 1987; West and Fenstermaker 1995). Gender ideologies adapted to meet various social needs (Martin 2004) and are a major component in structuring individual, interactional, and institutional power relations (Risman 2004). A dominant gender construct is sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). This along with other gender stereotypes affect the performance of real world applications, including online advertising (Sweeney 2013), toxic language detection (Park, Shin, and Fung 2018), machine translation and language technologies (Font and Costa-Jussà 2019; Dev et al. 2021; Vanmassenhove, Hardmeier, and Way 2018), search engines (Kopeinik et al. 2023; Kay, Matuszek, and Munson 2015; Albawardi and Jones 2023), and T2I systems (Bianchi et al. 2023; Naik and Nushi 2023).
Stereotypes & Amplification of Societal Biases in T2I Systems
Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone (Jain et al. 2021; Choi et al. 2020) demonstrating ways in which generative AI systems amplify extant social inequalities and patterns of over/under-representation. Bianchi and colleagues (2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts. They show using Stable Diffusion, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p. 4) and “a terrorist” generates brown faces with dark hair and beards. Qadri et al. (2023, p. 506) examine the cultural limitations of T2I systems by soliciting South Asian community experts to generate text prompts and annotate the resulting imagery, revealing T2I models “can participate in and scale such existing harmful media regimes of representation and amplify experiences of socio-cultural marginalization.” Similarly, Naik et al. (2023) explore exacerbation of social inequalities along axes of (perceived) gender, age, race, and geographical location, showing state-of-the-art T2I systems amplify real world inequities when compared to U.S. Bureau of Labor Statistics.
Figure 1: Examples of harm amplifictation for images generated with Stable Diffusion 2.1 with the input prompt shown underneath each image for sexually explicit content (A) and violent content (B). A gaussian blur or black box was applied on some images to limit exposure of harms to readers.
It is of note that stereotype amplification and harm amplification, although related, differ in key ways. Stereotype amplification in AI systems involves the reinforcement of stereotypes related to a group of people and does not necessarily pertain to safety. Conversely, harm amplification as we define it, occurs when AI systems exacerbates harms in the context of safety without necessarily targeting specific demographic groups. While stereotype amplification has been well studied, the formalization of harm amplification is novel. Furthermore, our understanding of harm amplification is related, but distinct, from extant work on bias amplification examining how patterns of bias appear in training data, and are amplified statistically in model outputs (Zhao et al. 2017; Wang and Russakovsky 2021; Zhao, Andrews, and Xiang 2023; Lloyd 2018; Taori and Hashimoto 2023; Mehrabi et al. 2021). In this context, bias amplification reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics” (Hall et al. 2022a, p. 1), often focusing on binary classification models (e.g., (Leino et al. 2019; Wang et al. 2022; Hall et al. 2022b)). While bias amplification typically addresses demographic disparities in classifier accuracy, harm amplification extends beyond predication accuracy and statistical disparities, aiming to understand how the generated model output’s safety relates to the user input.
Defining Harm Amplification
Our paper focuses on examining unsafe amplifications of harms in a specific set of generative AI - T2I systems. First, we define “unsafe” content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements. What could be considered “unsafe” or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations (Rismani et al. 2023; Dobbe 2022; Leveson 2020). For illustrative purposes in this paper, we explore two specific kinds of harmful representations to illustrate how harm amplification can be quantified in T2I systems, following the definitions of Hao et al. (2023):
  • Sexually explicit content: Depictions of explicit or graphic sexual acts, nudity (beyond human anatomy and presented in a sexually suggestive manner), sexualized body parts, or sex toys.
  • Violent content: Representations of realistic acts of violence, including but not limited to blood, dismemberment of body parts, and/or displays of aggression/force.
After harms and safety requirements have been established for a specific T2I system and context, we propose the following definition for harm amplification. Let $T$ be the space of all possible input prompts, $I$ be the space of all generated images, and $M$ be a text-to-image model that maps input prompts to generated images. Let $H$ be a harm function that measures the level of harm of the input or output of the model. For a given input prompt $t\in T$ and its corresponding generated image $i=M(t)$, we define harm amplification as occurring when:
$H(i)>H(t)+\tau$ (1)
where
  • $H(i)$ is the harm level of the output image alone
  • $H(t)$ is the harm level of the input prompt alone
  • $\tau$ represents a predefined threshold parameter
Thus, harm amplification can be defined as occurring when the output generated image is more harmful than the input text prompt. In this sense, amplification reflects when a T2I system increases the severity of the defined harms without being explicitly prompted to do so (e.g., sexually objectifying women in generated imagery). Thus, the notion of harm amplification we propose is concerned with understanding the relationships between the text input and resulting imagery. In this paper, we choose not to articulate a specific threshold for what constitutes as harm amplification as we believe this should be decided within a specific product or technology context with input from a wider range of stakeholders. Instead, we offer a methodological framework that demonstrates how to derive thresholds in various deployment scenarios and show how to employ this framework on an example dataset.
Framework to Identify Harm Amplification
In this section, we introduce three methods to quantify harm amplification. We then evaluate the efficacy of these methods on the human annotated Nibbler evaluation dataset (Quaye et al. 2024).
Methods
Quantifying harm amplification is a crucial component in assessing safety in T2I systems with respect to defined unsafe content. In this section, we introduce three statistical methods designed to measure harm amplification. Leveraging safety classifiers and embedding models, our approach aims to understand how the harms of the generated output images are amplified in the context of the user input prompts. Each of our methods has strengths and limitations that makes them well-suited for specific cases and available resources. This tailored approach ensures that our methods can be used across diverse contexts contributing to a more robust assessment of harm amplification.
Method 1 is most appropriate when researchers have access to a large amount of generated data and safety classifiers. In cases where data availability is limited, but there is access to safety classifiers, Method 2 may be preferable. Method 3 is best suited to low resource situations when researchers can only utilize open source models such as CLIP.
Method 1: Distribution-Based Thresholds for Discretized Text Harm Scores
The Distribution-Based Thresholds method utilizes: 1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and 2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image). Consider a set of text harm scores (${H(t):t\in T}$) and image harm scores (${H(i):i\in I}$). We first discretize the text harm scores into $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
$B_{j}=\{t\in T:l_{j}<H(t)\leq l_{j+1}\},\quad j=0,1,...,n-1$ (2)
For each $B_{j}$, the corresponding image harm scores $H(i):i=M(t),t\in B_{j}$ form a distribution $D_{j}$. Statistical measures, including mean ($\mu_{j}$) and standard deviation ($\sigma_{j}$), are calculated for each $D_{j}$. The non-fitted, raw image harm amplification threshold for each bucket is determined as $\mu_{j}+2\cdot\sigma_{j}$, or the 95th percentile, $P_{95,j}$ of the distribution. To provide a smooth representation of the calculated thresholds, a polynomial function can be fit through the raw threshold values across all buckets. In our experiments, we fit a first-degree polynomial with the corresponding equation to determine the image harm amplification threshold $HarmAmpThresh_{j}$ for each text harm bucket $B_{j}$:
$HarmAmpThresh(j)=b_{1}j+b_{0}$ (3)
where $b_{1}$ and $b_{0}$ are the coefficients of the fitted polynomial, determined by minimizing the least squares error:
$\min_{b_{1},b_{0}}\sum_{j=0}^{n-1}(P_{95,j}-(b_{1}j+b_{0}))^{2}$ (4)
Harm amplification is determined to occur for a pair $(t,i)$ if $t\in B_{j}$ and $H(i)>HarmAmpThresh(j)$. As mentioned previously, this method works best when there is access to large amounts of data to calculate statistical measures from. Additionally, this method performs well when text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier). Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
Method 2: Bucket Flip for Discretized Text and Image Scores
Utilizing text and image harm safety scores similar to Method 1, the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text. Consider a set of text harm scores (${H(t):t\in T}$) and image harm scores (${H(i):i\in I}$) obtained by safety classifiers. Assuming that the classifiers are well aligned, we can discretize text and image harm scores into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 2).
For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$. If the image harm bucket is higher than the text harm bucket ($B_{i}>B_{t}$), then we determine that harm amplification occurred.
Unlike Method 1, this method does not require any data and is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores. In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical and easy approach for identifying harm amplification.
Method 3: Image-Text Co-embedding Based Harm Scores
While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many researchers and practitioners may not have the resources to label such data and thus may not be able to train the necessary safety classifiers. In this case, existing large pre-trained image-text models can serve as a substitute. Large pre-trained image-text models, such as CLIP (Radford et al. 2021), embed images and their corresponding text labels such that the cosine distance between corresponding text and images is minimized in the embedding space. Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s CLIP embedding, $z_{i}$, is to a harmful concept in the embedding space than the prompt’s CLIP embedding, $z_{t}$, is. Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the sexually explicit harm category, we can use the words “sexual”, “porn”, etc. - see Appendix, Table 2), defining the texts as $\{h_{k}\}_{k=1}^{K}$ and their subsequent embeddings as $z_{h_{k}}$ These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space. We can define our harm function $H$ by calculating the average cosine similarity between a given text or image embedding and harm concepts:
$H(x)=\frac{1}{K}\sum_{k=1}^{K}cos\_sim(z_{h_{k}},z_{x})$ (5)
where $x$ can denote a text or an image and $cos\_sim(z_{h_{k}},z_{x})=\frac{\mathbf{z}_{h_{k}}\cdot\mathbf{z}_{x}}{\|\mathbf{z}_{h_{k}}\|\|\mathbf{z}_{x}\|}$. Harm amplification is then determined to occur for a pair $(t,i)$ if:
$H(i)-H(t)>\tau$ (6)
where $\tau>0$ is a predefined threshold. This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt. It is of note, that there may exist biases in the CLIP embedding space that may influence what is considered harm amplification. Applying this method using a debiased embedding space (Wang, Zhang, and Sang 2022) should be considered as a future direction.
Evaluation of Methods
Here, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset. These methods are then evaluated on an independent human annotated dataset where f1-scores, precision, and recall are derived.
Harm Amplification Measurement Method Sexually Explicit Violence
Precision Recall F1-Score Precision Recall F1-Score
Distribution-Based Thresholds 0.864 0.985 0.920 0.460 0.831 0.592
Bucket Flip 0.910 0.950 0.930 0.708 0.554 0.622
Image-Text Co-embedding 0.585 0.920 0.715 0.322 0.566 0.410
Table 1: Harm Amplification measurement methods evaluation on Nibbler data. Each method, when applied on the measurement dataset (Stable Diffusion 2.1 prompt, image dataset), yielded some criteria for determining whether harm amplification occurred for each harm type: sexually explicit and violence. These criteria were then applied to the Nibbler evaluation dataset where we had ground truth human annotation labels. Precision, recall, and F1-scores for sexually explicit content (left 3 columns) and violence (right 3 columns) were reported for each of the three measurement methods.
Method 1 Measurement Dataset
To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories. These prompts were sampled from dogfood user data aimed at testing generative AI models. By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment. While curated adversarial prompts can be valuable for specific safety assessments, our method aims to understand how harm amplification can occur over a large range of prompts (from benign to borderline to harmful). We used Stable Diffusion 2.1 (Rombach et al. 2022) to generate images for our prompt measurement dataset. Four images were generated per prompt and no NSFW or safety filters were applied during the image generation process. Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
Safety Classifiers and Image-Text Models
Pretrained safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2. Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1. For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al. 2021).
Evaluation Data & Metrics
To create the evaluation dataset, we sourced 1125 prompt, image pairs from the Adversarial Nibbler data challenge (Quaye et al. 2024). This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 (Open AI 2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) (Stability AI 2022). Five to six raters trained on AI safety policies annotated prompts and images separately for the presence of sexually explicit content or violent content. To derive a confidence score for sexually explicit and violence, we converted the ratings of the annotators into a proportional measure (i.e., if 3 out of 5 raters rated an image as violent, the violence score for that image would be 0.6 or 60%). The ground truth labels for harm amplification were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm). We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset. This evaluation was conducted separately for sexually explicit and violence harm amplification.
Evaluation of Methods in Detecting Harm Amplication
To quantify harm amplification using Method 1, Distribution-Based Threshold Method, we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence. Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix, Figure 5A). The 95th percentile score was then calculated for each distribution within its respective bucket with a first-degree polynomial being fit across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 5B) for each harm type. For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution. Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset. A precision of 0.864 and recall of 0.985 was observed when evaluating sexually explicit content in the Nibbler dataset (see Table 1). When evaluating violent harm amplification, we noted a precision of 0.460 and recall of 0.831.
Using Method 2, the Bucket Flip Method, we divided text and image harm scores obtained by safety classifiers applied on the Nibbler data into 5 even buckets. We then assessed for each prompt, image pair whether the image harm bucket was larger than the text harm bucket. This approach yielded better results than Method 1 with a precision of 0.910 and recall of 0.950 for sexually explicit harm amplification, and a precision of 0.708 and recall of 0.554 for violence harm amplification.
While we expect Method 3, Image-Text Co-embedding Method, to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources. We used CLIP to encode images and prompts, and defined sexually explicit harm concept using 15 words such as “porn”, “sexual”, and “nude” (see Appendix, Table 2). These concept words were sourced from internal adversarial testing and (Rando et al. 2022a) in their attempt to reverse-engineering Stable Diffusion’s NSFW filter. Adversarial testing was used to find 15 violence concept words such as “violence”, “weapons”, and “blood” (see Appendix, Table 2). Various harm amplification thresholds were applied to the evaluation dataset to obtain precision-recall curves (see Appendix, Figure 7). Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case. For these results we obtained a threshold that maximized the f1 score, and computed precision and recall from this number using the PR curves. As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.585, recall of 0.920) than violence harm amplification (precision 0.322, recall 0.566). Poorer performance on violence across methods could be attributed to less accurate classification models or noisier human labels. Future research should explore biases within classifiers and datasets and their impact on measuring harm amplification.
Examining Patterns of Harm Amplification
In the previous section, we showed how to apply different methodologies to measure harm amplification. Here, we explore patterns of harm amplification within the data, examining gender differences and applying explainability techniques to gain deeper insights of how different inputs affect harm amplification.
Gender Disparities
As described in the Related Works section, the amplification of social stereotypes disproportionately affects marginalized communities (Bianchi et al. 2023; Naik and Nushi 2023). In the evaluation dataset, we analyzed ground truth harm amplification rates and their disparate impacts on perceived genders. In this study, we recognize the dynamic and non-binary nature of gender, but ultimately chose to focus on perceived binary (male/female) gender expressions due to the ease in methodology. We also acknowledge that we cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
Annotations of perceived gender were collected for each image in the Nibbler evaluation dataset using gender classifiers (Baruah et al. 2022). We restricted our analysis to images in which there was a strict majority of a perceived gender. Images in which there were no faces or where the number of perceived males equaled that of perceved females were discarded. We then calculated the rates of harm amplification using the majority perceived gender. Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2),
Figure 2: Difference in harm amplification rates across perceived genders in the Nibbler evaluation dataset. Images containing females were significantly more oversexualized than males whereas there was no significant difference in harm amplification of violence. $*p<0.05,\ **p<0.01,\ ***p<0.001$
in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011). There was no significant difference in violence harm amplification across perceived genders ($p=0.492$). We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
Applying Explainability Techniques
To gain deeper insights into the mechanisms of harm amplification, we employed explainability techniques such as attribution maps and counterfactual analysis. We utilized Diffusion Attentive Attribution Maps (DAAM) (Tang et al. 2023) to visualize the relationship between input words and output image pixels, leveraging cross attention maps in diffusion models. Figure 3 shows attribution maps for two examples in our dataset that exhibited notable harm amplification.
Figure 3: Attribution maps (Tang et al. 2023) for notable harm amplification prompts.
For the prompt “Person venting frustration.”, the attribution maps show that the words “frustration” and to a lesser extent “venting” contribute to the blood splatter observed in the image. In contrast, for the prompt “black gay man”, while the word “gay” seems to be attributed to some facial and neck features, the attribution maps alone do not capture what aspects of the prompt contribute the man being depicted shirtless.
To further investigate cases where attribution maps provide limited insights, we ran a counterfactual analysis. For the prompt “black gay man”, we generated counterfactual variations by altering race and sexual orientation terms. Using the same generation seed, we produced images from these counterfactual prompts (Figure 4).
Our analysis revealed that prompts altering the race term still exhibited similar levels of harm amplification. However, when modifying the sexual orientation term, we observed a significant reduction in harm amplification suggesting that the sexual orientation term played a more substantial role in triggering sexually explicit harm amplification than the race term. These findings show the complex interaction between different demographic descriptors and harm amplification in T2I models and highlight the need for future research of intersectional identities and harm amplification.
Figure 4: Counterfactual analysis for the prompt “black gay man”, altering the sexual orientation term (red) and the race term (blue).
Discussion
In this paper, we provide the first formalization of harm amplification in T2I models, defining it as occurring when the generated output image reflects more harmful representations that were not explicit in the text input. Using two specific kinds of harmful representations (sexually explicit content and violent content), we provided a framework of different methods to identify instances of harm amplification under different resource and deployment scenarios. We then evaluated this framework on an independent human annotated dataset.
Reduced Reliance on Human Annotation
Our framework significantly reduces the reliance on human annotations. By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets. While human annotations are valuable for evaluating the method’s efficacy on held out datasets, they are not required for the core measurement process. Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive. It is of note that machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al. 2021) — can also reflect social biases. These biases may stem from the subjective nature of defining ground truth in safety. Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
Understanding Demographic Disparities in Harm Amplification
In analyzing rates of harm amplification in the human annotated evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content. Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images. These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender. Furthermore, we show through explainability analysis that sexual orientation also affects the severity of harm amplification. Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts on different sociodemographic groups.
Future Work
While our work lays a foundation to better understand and quantify harm amplification in T2I systems, further research is needed to more holistically understand the phenomenon and develop strategies for its reduction in T2I models. One future research potential includes an in-depth exploration of the role of training data and whether biases in training data lead to the reinforcement of stereotypes for certain groups as we observed in our results pertaining to the oversexualization of women. This work will be pivotal to understanding the root cause of harm amplification, thereby guiding the development of solutions to reduce its impact in T2I models. Furthermore, there is a need for the expansion of T2I safety mitigation techniques, similar to Safe Latent Diffusion (Schramowski et al. 2023), to reduce harm amplification as these strategies are necessary in minimizing the exposure of unwanted harms to users. The development of these mitigation techniques will contribute to the overall responsible deployment of T2I systems, supporting the increased protection and safety of users.
Acknowledgements
We thank Kathy Meier-Hellstern, Alexander Vorontsov, Vinodkumar Prabhakaran, and Atoosa Kasirzadeh for their feedback and suggestions. We also thank Auriel Wright, Candice Schumann, Lora Aroyo, and Alicia Parrish for discussions on analysis and writing.
Appendix A Appendix
Figure 5: Method 1: Distribution-based thresholds applied on the measurements dataset. A) Machine annotations for sexually explicit harm scores were obtained for text and images. Text sexually explicit scores were bucketed into 5 buckets. Distributions of the image sexually explicit scores were then derived for images with corresponding text in that bucket. 2 standard deviations above the mean was used as an initial raw threshold where we then fitted a 1 degree polynomial to obtain new fitted thresholds (right most image). B) We then repeated the same procedure for machine annotations for violence.
Sexually Explicit
porn sexual
nude sex
18+ naked
nsfw dick
vagina explicit content
uncensored fuck
nipples naked person
sexy
Violence
violence gore
blood attack
bloodshed war
horror fight
weapons injury
death pain
wound brutality
harm
Table 2: Harm concepts for Method 3: Image-Text Co-embedding. These harm concepts were meant to approximate the embedding of sexually explicit and violence.
Figure 6: Confusion matrix for each method (left: Distribution-Based Thresholds, middle: Bucket Flip, right: Image-Text Co-embedding) evaluated on the Nibbler dataset.
Figure 7: Precision-recall curves for the Image-Text Co-embedding method for sexually explicit content (left) and violent content (right).
Harm Amplification Measurement Method Female Male
Precision Recall F1-Score Precision Recall F1-Score
Distribution-Based Thresholds 0.858 0.979 0.915 0.792 0.987 0.878
Bucket Flip 0.913 0.932 0.923 0.860 0.961 0.908
Image-Text Co-embedding 0.656 0.911 0.763 0.308 0.844 0.451
Table 3: Efficacy of the proposed methods across the two perceived gender expressions for sexually explicit harm amplification.